Documenta Mathematica

Journal der Deutschen Mathematiker-Vereinigung

Band 3, 1998

Peter Heinzner The Minimum Principle

from a Hamiltonian Point of View  1–14

Stefan Kebekus Simple Models

of Quasihomogeneous Projective 3-Folds  15–26

Annette Huber, Jörg Wildeshaus Classical Motivic Polylogarithm

According to Beilinson and Deligne  27–133

Vladimir Chernousov An Alternative Proof of Scheiderer’s Theorem on the Hasse Principle

for Principal Homogeneous Spaces  135–148

Meike Schäffler Partition Regular Systems

of Linear Inequalities  149–187

Detlev W. Hoffmann and Jean-Pierre Tignol On 14-dimensional quadratic forms in I^{3}, 8-dimensional forms in I^{2},

and the common value property  189–214

Marc A. Rieffel

Metrics on States from Actions of Compact Groups  215–230

Bernhard Keller On the Cyclic Homology

of Ringed Spaces and Schemes  231–259

Michael Scheutzow, Heinrich v. Weizsäcker Which Moments of a Logarithmic Derivative

Imply Quasiinvariance?  261–272

Haruzo Hida

Global Quadratic Units and Hecke Algebras  273–284

Erez M. Lapid

A note on the Global Langlands Conjecture  285–296

Annette Huber, Jörg Wildeshaus Correction to the Paper ”Classical Motivic Polylogarithm

According to Beilinson and Deligne”  297–299

Annette Werner Local Heights on Abelian Varieties

and Rigid Analytic Uniformization  301–319

Mina Teicher

Chern Classes of Fibered Products of Surfaces  321–332

M. S. Abual-Rub Long Range Diffusion Reaction Model

on Population Dynamics  333–340

Oleg T. Izhboldin

Motivic Equivalence of Quadratic Forms  341–351

Volker Bach, Jean-Marie Barbaroux, Bernard Helffer, Heinz Siedentop Stability of Matter for the Hartree-Fock Functional

of the Relativistic Electron-Positron Field  353–364